OpenStudy (anonymous):

18000= 30(2^12x) anyone?

OpenStudy (anonymous):

divide by 30 \[600=2^{12x}\]change of base formula \[12x=\frac{\ln(600)}{\ln(2)}\]divide by 12 \[x=\frac{\ln(600)}{12\ln(2)}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw